# Deep Modularity Networks with Diversity–Preserving Regularization

### v1

@@ 5 Results
DMoN 11.00 72.39 43.65 46.24 4.92 75.72 30.26 42.17 8.66 57.03 22.68 45.18
DMoN–DPR(D) 10.23 72.97 45.31 50.00 4.75 75.64 30.61 42.76 8.66 57.02 22.68 45.19
DMoN–DPR(V) 11.64 70.80 44.27 48.18 4.93 75.59 30.46 42.26 8.66 57.02 22.68 45.18
DMoN–DPR(E) 11.00 72.37 43.63 46.25 5.02 75.43 30.52 42.54 8.67 57.02 22.68 45.19
DMoN–DPR(DV) 10.37 72.68 45.41 50.13 4.76 75.51 30.59 42.85 8.66 57.02 22.68 45.19
DMoN–DPR(DE) 10.33 72.88 45.18 49.88 4.86 75.35 30.77 43.27 8.66 57.02 22.67 45.19
DMoN–DPR(VE) 11.64 70.77 44.15 48.09 5.09 75.26 30.86 42.86 8.66 57.01 22.68 45.18
DMoN–DPR(DVE) 10.41 72.65 45.43 50.06 4.85 75.14 30.77 43.42 8.66 57.02 22.67 45.18
@@
Method Coauthor CS Coauthor Phys
@@
DMoN 18.80 72.43 69.45 58.83 13.81 63.35 52.40 46.80
DMoN–DPR(D) 19.01 72.17 70.56 61.09 16.45 56.93 53.99 50.96
DMoN–DPR(V) 18.89 72.41 69.76 59.10 13.45 59.38 54.03 51.89
DMoN–DPR(E) 19.73 70.97 71.34 60.40 13.80 60.04 53.17 50.05
DMoN–DPR(DV) 18.95 72.26 71.33 61.81 13.46 56.85 54.09 56.36
DMoN–DPR(DE) 19.88 70.56 70.84 61.76 14.22 56.10 54.71 56.94
DMoN–DPR(VE) 19.70 71.01 71.45 60.56 12.85 55.72 50.14 54.52
DMoN–DPR(DVE) 19.74 70.68 71.08 62.41 13.46 55.27 54.31 58.86
5.1 Effectiveness of DMoN–DPR in Clustering
To assess the effectiveness of DMoN–DPR, we compare the mean values of conductance ($C$), modularity ($Q$), NMI, and pairwise F1 measure across 10 seeds achieved with DMoN–DPR to those obtained with DMoN, as done in (Tsitsulin et al., 2023) (full results can be found in Section B). Tables 2 (Cora, CiteSeer, PubMed) and 3 (Coauthor CS, Coauthor Physics) summarize the results, where graph metrics (C, Q) do not rely on ground-truth labels, whereas ‘Labels’ metrics (NMI, F1) compare to known class labels.
@@ 5.1 Effectiveness of DMoN–DPR in Clustering
On Cora, where the diversity of node features and class labels is relatively high, incorporating the distance-based term (D) consistently enhances performance over the baseline DMoN, especially on label-based metrics (NMI and F1). The combinations that include the distance term—such as DMoN–DPR(DV), DMoN–DPR(DE), and DMoN–DPR(DVE)—not only improve cluster separability but also boost label-based performance, with DMoN–DPR(DV) and DMoN–DPR(DVE) achieving some of the best F1 and NMI scores. The entropy (E) and variance (V) terms alone are less impactful, but when paired with distance, they synergize to produce more balanced and diverse clusters. On CiteSeer, which also has a reasonably complex feature space, the introduction of these regularizations yields moderate gains, with combinations that include the entropy term slightly improving label-based metrics. In contrast, PubMed, which has a less diverse feature space, sees only marginal benefits. Here, the baseline DMoN already performs well, and the added terms provide limited additional advantage.
@@
We observe a similar pattern on the two Coauthor datasets. Both of these datasets have richer feature spaces than Cora, CiteSeer, and PubMed; however, Coauthor Physics stands out with the largest number of features among all the datasets under consideration (8,415). As shown in Table 3, this higher dimensionality helps amplify the benefits of our diversity-preserving regularizations, especially those involving the distance-based term (D). Notably, on Coauthor Physics, the pairwise F1 score jumps from 46.80 (DMoN) to 58.86 (DMoN–DPR(DVE)), indicating that combining distance, variance, and entropy can substantially enhance the separability of clusters and the alignment with true labels.
On Coauthor CS, our proposed regularizations also offer improvements over baseline DMoN, most evident when distance (D) is paired with either variance (V) or entropy (E). For example, F1 increases from 58.83 (DMoN) to 62.41 (DMoN–DPR(DVE)). Although these gains are meaningful, they are somewhat more moderate than those on Coauthor Physics—likely due to differing distributions of class labels and feature dimensions. Overall, these findings suggest that introducing distance-, variance-, and entropy-based regularizations becomes increasingly beneficial as the dimensionality and diversity of node features grow, leading to more pronounced performance boosts on datasets such as Coauthor Physics.
Analyzing the Influence of Regularization on Clustering Performance.
Our experiments show that the impact of the distance-based, variance-based, and entropy-based regularizations depends heavily on both the feature richness and label diversity of each dataset. On Coauthor CS, and Coauthor Physics, which have relatively higher-dimensional or more varied feature spaces, we observe notable gains in NMI and F1 upon adding these diversity-promoting terms—especially the distance-based (D) regularization, which strongly boosts inter-cluster separation. For instance, DMoN–DPR(D) and its variants (e.g., DMoN–DPR(DV), DMoN–DPR(DE), DMoN–DPR(DVE)) yield significant improvements in label alignment on Coauthor Physics, where the F1 score rises by over 10 percentage points compared to baseline DMoN.
Meanwhile, the variance-based (V) term is helpful in ensuring clusters capture a wider range of node features, an effect that is especially apparent on datasets like Coauthor Physics. Although the single V-only variant often yields moderate gains on datasets such as Cora, it generally performs best when combined with the distance or entropy regularizers. The entropy-based (E) term aims to produce more balanced cluster distributions, proving modestly beneficial on datasets like CiteSeer and effective on Coauthor CS and Coauthor Physics; however, on PubMed, which has lower feature dimensionality, we see little to no improvement (the numbers remain nearly unchanged from DMoN). This indicates that entropy-based balancing may not substantially help in more homogeneous datasets.
Moreover, combining all three terms (DMoN–DPR(DVE)) often yields the highest F1 scores across multiple datasets (e.g., a leap from 58.83 to 62.41 on Coauthor CS and 46.80 to 58.86 on Coauthor Physics). These results suggest that multiple regularizers can work in synergy, but the exact benefits depend on a dataset’s inherent complexity. Overall, the findings highlight the importance of tailoring regularization strategies to both the graph structure and feature diversity of the data to achieve the most effective clustering outcomes.
Trade-Offs Between Structural and Label-Based Metrics
Our results highlight a trade-off between metrics that prioritize graph structure (e.g., Modularity ($Q$) and Conductance ($C$)) and those that evaluate label alignment (e.g., NMI and F1). This trade-off emerges because diversity-preserving regularization terms influence both the feature and label spaces of clusters, leading to distinct outcomes depending on the dataset characteristics and application needs.
For datasets with rich feature spaces (e.g., Coauthor CS and Coauthor Physics), the diversity-enhancing terms significantly boost label-based metrics like NMI and F1. These gains stem from the formation of more homogeneous clusters that align closely with ground-truth labels. However, this improvement in label alignment can come at the expense of structural metrics such as modularity, which reward densely connected clusters in the graph topology. For instance, while DMoN–DPR(DVE) achieves the highest F1 scores on Coauthor Physics, it exhibits a decrease in modularity compared to baseline DMoN.
Conversely, on datasets with less diverse or lower-dimensional features (e.g., PubMed), the structural metrics are less affected by the introduction of diversity-preserving regularization terms, and the benefits to label-based metrics are more modest. In such cases, the original DMoN, which focuses on structural cohesion, remains a competitive choice for purely unsupervised clustering tasks.
Ultimately, the decision to prioritize label alignment or structural metrics should be informed by the application context. For tasks involving semi-supervised learning or node classification, where accurate label recovery is essential, DMoN–DPR offers clear advantages. In contrast, for community detection or unsupervised settings where graph structure is paramount, the baseline DMoN may be preferable. This balance between competing objectives underscores the need to tailor clustering approaches to the specific demands of the dataset and task at hand.
5.2 Visualization of Clusters
The clustering results visualized in Figure 1 demonstrate that DMoN–DPR generally outperforms DMoN across the benchmark datasets Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics. For Cora, DMoN–DPR produces tighter, more well-separated clusters with improved alignment to the ground truth labels, contrasting with the more scattered and less defined clusters formed by DMoN. Similarly, for CiteSeer, DMoN–DPR exhibits better cluster compactness and reduced overlap, though the gains are somewhat less pronounced than in Cora. On the other hand, PubMed remains challenging for both methods, resulting in notable overlap and scattered clusters; yet, DMoN–DPR still provides slightly tighter groupings. Across these three datasets, the diversity-preserving regularization in DMoN–DPR effectively enhances both inter-cluster distinctiveness and intra-cluster variability, producing more meaningful and interpretable partitions. Turning to the Coauthor CS and Coauthor Physics datasets, a similar trend emerges. While DMoN adequately captures overarching modular structures, it can occasionally struggle to separate nuanced subgroups. In contrast, DMoN–DPR consistently yields more cohesive and better-separated clusters, offering clearer decision boundaries and higher alignment with the true labels. This improvement is particularly notable in Coauthor CS, where DMoN–DPR visibly reduces overlap and sharpens cluster delineations. Although the inherently complex Coauthor Physics dataset can induce some overlap in both methods, DMoN–DPR still produces more robust clusters overall. These findings underscore the effectiveness of promoting diversity within hidden representations, as DMoN–DPR not only excels in forming compact and accurate clusters for datasets with greater feature variability but also remains competitive on more challenging datasets.
@@ 5.3 Execution Times and Runtime Analysis
Dataset DMoN DMoN–DPR
Cora 32 35
CiteSeer 76 71
PubMed 310 300
Coauthor CS 722 751
Coauthor Physics 2107 1946
@@
Furthermore, running on CPU can introduce variability from scheduling, caching, and background processes, so small differences in the recorded times may not be statistically meaningful. Thus, although DMoN–DPR performs slightly more computations, the extra cost is overshadowed by the main bottleneck operations, resulting in comparable or even lower measured runtimes in practice.
@@ 5.4 Limitations and Future Directions
Despite the improvements introduced by our diversity–preserving regularization, DMoN–DPR has a few notable limitations. First, tuning the regularization weights for distance, variance, and entropy remains a manual process. Determining the best values for these hyperparameters can be time-consuming and sensitive to each dataset’s characteristics. An interesting direction for future work is to develop an automated or adaptive scheme that learns these weights during training, making the method more robust and less reliant on exhaustive search.
Second, while DMoN–DPR yields substantial gains on datasets with a large or diverse feature space, the improvements are often marginal for datasets with lower feature dimensionality (e.g., PubMed). Nonetheless, our visualization in Figure 2 suggests that even in these cases, DMoN–DPR tends to produce slightly more coherent cluster structures compared to the baseline, reflecting the benefit of promoting diversity even when numerical metrics (e.g., NMI, F1) show limited change. Future research could investigate whether alternative or more specialized regularizers can drive greater numerical improvements on lower-dimensional datasets.

### v2

@@ 5 Results
DiffPool 15.99 62.78 40.13 46.55 7.92 66.69 33.40 47.83
MinCut 13.65 71.79 37.74 39.15 6.19 75.21 25.21 35.28 11.14 54.66 22.48 41.31
DMoN 10.53 72.77 43.92 46.93 4.86 75.45 29.67 42.46 8.61 57.13 22.39 43.23
DPR(D) 10.95 72.20 43.98 47.51 4.95 75.56 30.09 42.46 8.60 57.14 22.38 43.24
DPR(V) 11.87 70.54 43.34 46.39 4.95 75.17 29.76 42.85 8.61 57.13 22.39 43.23
DPR(E) 10.49 72.77 44.00 47.02 4.93 75.25 29.95 42.96 8.62 57.11 22.36 43.22
DPR(DV) 10.89 72.07 44.40 47.36 5.09 75.18 30.50 43.25 8.60 57.14 22.38 43.24
DPR(DE) 10.88 72.30 44.30 47.53 5.07 75.18 30.47 43.13 8.60 57.12 22.38 43.25
DPR(VE) 11.86 70.51 43.24 46.36 5.05 75.01 30.17 43.27 8.60 57.13 22.37 43.21
DPR(DVE) 10.93 72.05 44.37 47.32 5.16 75.00 30.25 43.00 8.61 57.11 22.36 43.24
@@
Method Coauthor CS Coauthor Physics
@@
DiffPool 18.19 62.24 53.92 53.47 13.91 57.44 56.22 51.71
MinCut 21.36 71.58 64.33 49.00 13.93 61.75 51.39 47.79
DMoN 18.63 72.60 69.26 59.26 13.70 63.45 53.50 47.51
DPR(D) 18.89 72.30 71.17 61.82 16.44 56.89 53.49 50.99
DPR(V) 18.75 72.50 69.56 59.94 13.18 58.59 53.93 52.78
DPR(E) 19.85 70.92 71.58 61.33 13.49 59.33 52.83 51.09
DPR(DV) 19.14 71.96 70.72 61.35 13.43 56.48 55.84 57.99
DPR(DE) 19.81 70.57 70.96 62.36 14.17 56.89 54.02 55.02
DPR(VE) 19.89 70.81 71.47 61.33 13.11 55.88 49.95 53.58
DPR(DVE) 19.69 70.71 71.28 62.67 12.84 55.47 53.50 57.96
Table 3: t-statistics and p-values for all datasets (rounded to two decimal places).
Metric Cora CiteSeer PubMed Coauthor CS Coauthor Physics
t-stat p-val t-stat p-val t-stat p-val t-stat p-val t-stat p-val
Conductance -0.88 0.40 -1.62 0.14 0.37 0.72 -5.25 $1e^{-6}$ 0.51 0.62
Modularity 1.81 0.10 1.17 0.27 0.21 0.84 5.88 $1e^{-6}$ 17.56 $1e^{-6}$
NMI -0.63 0.54 -1.56 0.15 0.23 0.83 -4.64 $1e^{-6}$ -2.49 0.03
F1 -0.40 0.70 -1.28 0.23 -0.95 0.37 -2.51 0.03 -14.65 $1e^{-6}$
5.1 Effectiveness of DMoN-DPR in Clustering
To assess the effectiveness of DMoN-DPR, we compare the mean values of conductance ($C$), modularity ($Q$), NMI, and pairwise F1 measure across 10 randomly selected seeds achieved with DMoN-DPR (denoted as DPR in all the tables) to those obtained with our baselines. Tables 1 (Cora, CiteSeer, PubMed) and 2 (Coauthor CS, Coauthor Physics) summarize the results, where graph metrics (C, Q) do not rely on ground-truth labels, whereas ‘Labels’ metrics (NMI, F1) compare to known class labels.
For completeness, we defer supporting analyses to the appendix: (i) quantitative diagnostics of feature-space diversity (average/min inter-centroid distance, average intra-cluster variance, Silhouette score) (Appendix B), (ii) t–SNE visualizations of the learned clusters versus ground-truth labels (Appendix C), (iii) ablation and hyperparameter-selection studies that isolate the effect of each DPR term (Appendix D), and (iv) per-seed scores and full tables with error bars (Appendix E).
@@ 5.1 Effectiveness of DMoN-DPR in Clustering
Overall, the results confirm that adding diversity-preserving regularizers to DMoN almost never harms the purely topological scores—conductance ($C$ ↓) and modularity ($Q$ ↑)—while consistently lifting label-aware metrics such as NMI and F1. On Cora, DPR(DV) achieves the best NMI (44.40%) and only narrowly trails DPR(DE) on F1. On CiteSeer, although DiffPool attains the highest NMI (33.40%) and F1 (47.83%), DPR variants remain competitive without the instability that causes DiffPool to collapse on PubMed  [^1] With a single DiffPool layer and our fixed cluster budget ($C=3$), the link-prediction loss on the 19 k-node PubMed graph stagnated and the assignment matrix collapsed to one cluster. Increasing the budget restores numerical stability, but yields only marginal label quality and would also violate our fixed-$C$ protocol, so we omit DiffPool scores for PubMed.. Among the DMoN-family methods, DPR(DV) is the strongest on CiteSeer. On PubMed, all DPR variants remain stable, in contrast to DiffPool, while MinCut yields the highest NMI (22.48%) and DPR(DE) edges out others in F1. The fact that DPR(D) and DPR(DV) match the best conductance (8.60%) confirms that the regularizers maintain cut quality.
@@
On the more feature-rich [^2] We define feature-rich graphs as those that exhibit (i) high feature dimensionality and (ii) high average Shannon entropy per node. Please see Appendix A for more detail. Coauthor graphs, diversity-preserving regularization brings substantial gains. On Coauthor CS, DPR(DVE) achieves a new best in F1 (62.67%) while DPR(E) sets the highest NMI (71.58%), both outperforming the already strong DMoN baseline without sacrificing conductance or modularity. On Coauthor Physics, DPR(DVE) obtains the lowest conductance (12.84%) and, along with DPR(DV), surpasses 57% F1—demonstrating that the full three-term objective generalizes well to larger, noisier graphs.
Key Insights.
Therefore, our results show that the effects of distance-based (D), variance-based (V), and entropy-based (E) regularizations depend on dataset feature richness. On feature-rich datasets like Coauthor CS and Physics, adding these terms—especially D—significantly boosts NMI and F1, with DMoN-DPR(DV) raising F1 by over 10 percentage points. The V term increases assignment dispersion across nodes and performs best when combined with D. The E term, used with a small positive weight, gradually sharpens per-node assignments, which can indirectly improve cluster balance during early exploration on feature-rich datasets, but offers little benefit on simpler datasets like PubMed; the collapse regularizer remains the primary mechanism for cluster-size balance. Overall, combining D and V often yields the highest performance, underscoring the need to tailor regularization to dataset characteristics for optimal clustering.
Statistical Significance.
To strengthen our claims, we conduct a paired t‐test on Conductance, Modularity, NMI, and F1, comparing DMoN vs. the best‐performing DMoN–DPR variants. We ensure both methods use exactly the same seeds in each dataset. The results are summarized in Table 3. On the citation benchmarks we adopt the strongest variant for each graph (DV on Cora and CiteSeer, DVE on PubMed). Although the average NMI and F1 of DMoN–DPR surpass or match the vanilla DMoN baseline, the paired two-tailed $t$-tests in Table 3 show p-values above 0.10. This indicates that, given the modest sample of random seeds, the improvements are encouraging but not yet conclusive for these sparsely featured graphs. The picture is markedly different on the feature-rich coauthor networks. The gains on NMI and F1 metrics are highly significant, indicated by $p\leq 0.05$, as seen in Table 3. These results validate our hypothesis that encouraging latent cluster dispersion is most beneficial when node attributes are abundant and heterogeneous, and they confirm that DMoN–DPR maintains the stability advantage of the original objective while translating it into measurably better community recovery.
Trade-Offs Between Structural and Label-Based Metrics.
Our results reveal a clear trade-off between structural metrics (e.g., Modularity, Conductance) and label-based metrics (e.g., NMI, F1). Diversity-preserving regularization improves label alignment, especially on feature-rich datasets like Coauthor CS and Physics, by forming clusters that better match ground-truth labels—though sometimes at the cost of structural cohesion (e.g., lower modularity). Conversely, on datasets with less diverse or lower-dimensional features (e.g., PubMed), the structural metrics are less affected by the introduction of diversity-preserving regularization terms, and the benefits to label-based metrics are more modest. As a result, DMoN-DPR emerges as a more effective choice for supervised tasks prioritizing NMI and F1, whereas vanilla DMoN remains competitive for unsupervised scenarios where conductance and modularity are paramount. Tailoring the clustering strategy to the dataset and task requirements is therefore essential.
@@ 5.2 Execution Times and Runtime Analysis
Method Cora CiteSeer PubMed Coauthor CS Coauthor Physics
DMoN 32 76 310 722 2107
DMoN–DPR 35 71 300 751 1946
@@ 5.3 Limitations and Future Directions
While DMoN-DPR introduces clear benefits, it also presents a few limitations. First, tuning the diversity-preserving weights for distance, variance, and entropy remains manual. Though our empirical studies revealed consistent patterns: high-dimensional datasets like Coauthor CS and Physics benefited from larger distance weights (1 or 10), variance weights of 1 or 0.1 worked well across most datasets, and entropy value of 0.1 served as a good upper bound. A promising direction for future work is to develop an adaptive scheme that learns these weights during training, reducing reliance on manual tuning. Second, the gains from diversity regularization are more pronounced on feature-rich datasets. On lower-dimensional datasets like PubMed, improvements in NMI and F1 are modest—though visualizations (Figure 1) suggest that DMoN–DPR still forms more coherent clusters than the baseline. Future research could explore alternative regularizers to better capture structure in such settings.